# Improved Counting under Continual Observation with Pure Differential Privacy

### v1

@@ Abstract
Counting under continual observation is a well-studied problem in the area of differential privacy. Given a stream of updates $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\in\{0,1\}$ the problem is to continuously release estimates of the prefix sums $\sum_{i=1}^{t}\mathbf{x}_{i}$ for $t=1,\dots,T$ while protecting each input $\mathbf{x}_{i}$ in the stream with differential privacy. Recently, significant leaps have been made in our understanding of this problem under approximate differential privacy, aka. $(\varepsilon,\delta)$-differential privacy. However, for the classical case of $\varepsilon$-differential privacy, we are not aware of any improvement in mean squared error since the work of Honaker (TPDP 2015). In this paper we present such an improvement, reducing the mean squared error by a factor of about 4, asymptotically. The key technique is a new generalization of the binary tree mechanism that uses a $k$-ary number system with negative digits to improve the privacy-accuracy trade-off. Our mechanism improves the mean squared error over all “optimal” $(\varepsilon,\delta)$-differentially private factorization mechanisms based on Gaussian noise whenever $\delta$ is sufficiently small. Specifically, using $k=19$ we get an asymptotic improvement over the bound given in the work by Henzinger, Upadhyay and Upadhyay (SODA 2023) when $\delta=O(T^{-0.92})$.
@@ 1 Introduction
Private data analysis describes the setting in which we want to facilitate queries on datasets while safeguarding the privacy of the individuals making up the dataset. One increasingly popular framework for this area of research is referred to as differential privacy [12], which gives exact probabilistic guarantees for individuals making up such datasets, provided that algorithms fulfilling the corresponding definitions are used in the analysis. The trade-off between the utility and the privacy guarantees derived from these algorithms is heavily dependent on the exact nature of the queries. One especially challenging setting is that of continual observation [13, 15], in which the dataset arrives one entry at a time, and where we immediately after each entry arrives are expected to answer a query based on the entire dataset received up to that point in time. The challenge derives from the fact that once an element has arrived, it may influence all future queries.
One of the founding problems in this area is referred to as counting under continual observation, or continual counting for short [13, 6]. It describes the following problem: a binary stream $\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\dots\in\{0,1\}$ arrives one item at a time, where $\mathbf{x}_{t}$ is received at time $t$, and we are then asked to output the number of $1$s seen thus far. We consider two input streams $\mathbf{x},\mathbf{x}^{\prime}$ as neighbouring if they differ at exactly one index, and study the design of differentially private algorithms for continual counting under this neighbouring relation. While we will focus on the case when there is a known upper bound $T$ on the length of the stream ($\mathbf{x}\in\{0,1\}^{T}$), our results can be extended to the unbounded case using a standard reduction in [6].
Computing prefix sums on binary streams may appear as a toy problem but we note that: 1) it is a fundamental problem still with gaps between lower and upper bounds [13, 19, 16, 14, 9], and 2) it is used as a subroutine in various applications ranging from private learning [24, 22, 8, 11, 7] to histogram estimation [4, 5, 21, 26].
@@ 1.1 Our contributions
Our main contribution in this paper is an easy-to-implement tree-aggregation-based mechanism for continual counting under $\varepsilon$-DP, with properties stated in Theorem 1.1.
@@
Here and elsewhere $\log_{k}$ is the base $k$ logarithm, and $\log(\cdot)=\log_{2}(\cdot)$. To the best of our knowledge, this is the smallest mean squared error known for any $\varepsilon$-DP mechanism, up to lower-order asymptotically vanishing terms. It improves the error achieved by applying techniques in [20] to the binary tree mechanism (as implemented in [22] and referred to elsewhere as ?Honaker Online?) by a factor 4 when $k=19$. While this is interesting in its own right, we note that this improvement in the leading constant has implications on how the error compares to $(\varepsilon,\delta)$-DP mechanisms. For $\delta=O(T^{-0.92})$, our bound on the error beats the bound provided in [19] for the ?optimal? $(\varepsilon,\delta)$-DP matrix mechanism using Gaussian noise. This phenomenon is discussed in Section 5, but in a nutshell a sufficiently small $\delta$ must cause the error to greatly increase if the mechanism is based on Gaussian noise.
@@ 1.2 Overview of technical ideas
We start out by building a complete $k$-ary tree of height $h=\lceil\log_{k}(T+1)\rceil$ on top of the input $\mathbf{x}\in\{0,1\}^{T}$, where the first $T$ leaves stores $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$ and each internal vertex stores the sum of the leaves below it. We refer to the level $\ell$ of the leaves to be zero. We privately release the entire tree storing partial sums on $\mathbf{x}$, barring the root. To analyze this, it turns out that expressing a time step $t$ as a $k$-ary number, i.e., as $h$ digits in $\{0,\dots,k-1\}$, i.e. $\mathbf{t}\in\{0,\dots,k-1\}^{h}$, gives a direct means of computing which vertices contribute to a given prefix sum. With the convention that $\mathbf{t}_{1}$ is the least-significant digit, $\mathbf{t}_{\ell}\leq k-1$ then describes the number of left-children vertices on level $\ell$ that are added together to produce an estimate for the prefix sum up to $t$.
Naturally, higher arity trees are chosen due to their ability to accommodate a larger number of leaves with a smaller height compared to lower arity trees. However, this advantage then comes at the cost of potentially requiring more vertices per level (up to $k-1$) to generate a desired prefix sum estimate. To clarify the weaknesses of this setting consider the example illustrated in Figure 1, where a $5$-ary tree is depicted. It can be observed that a tree with a height of $2$ is sufficient to compute the prefix sum for $t=24$. However, to achieve this, it is necessary to add up to $4$ vertices from each level of the tree.
@@
To reduce the number of vertices added together per level, we introduce subtraction. The main idea is that on each level $\ell$, whenever we are adding $\geq\lceil\frac{k}{2}\rceil$ left-children vertices, we can instead add their parent on level $\ell+1$, and subtract the $\leq\lfloor\frac{k}{2}\rfloor$ right-children. To analyze the structure of the resulting prefix sums when subtraction is introduced, it turns out that rather than expressing $t$ in digits from 00 to $k-1$, we should instead allow for negative digits. Focusing on the case of odd $k$, we thus represent $t$ as $\mathbf{t}\in\{-\frac{k-1}{2},\dots,\frac{k-1}{2}\}^{h}$. The interpretation of $\mathbf{t}$ becomes: if $\mathbf{t}_{\ell}\geq 0$, add $\mathbf{t}_{\ell}$ left-children on level $\ell$, otherwise, subtract $|\mathbf{t}_{\ell}|$ right-children on level $\ell$. Identifying that $\left\lVert\mathbf{t}\right\rVert_{1}$ is equal to the number of vertices used to produce the estimate at time $t$, we can analyze the average number of vertices added for an estimate from this representation by studying $\mathbf{t}$ for $t\in[1,T]$. This gives roughly, a factor-2 reduction in the mean squared error versus not using subtraction.
An example of our improvement, for a single query at $t=8$, is illustrated in Figure 2, which is a ternary tree for $T=9$. In this case, a height of $2$ is sufficient and by using only addition, we would need to use a minimum of $4$ vertices. However, if subtraction is allowed we can compute the prefix sum using only $2$ vertices.
@@ 1.3 Related work
Historically, the first mechanisms with good error are typically credited to Chan, Shi, and Song [6] and Dwork, Naor, Pitassi, and Rothblum [13], though similar constructions were proposed independently by others [18, 17]. These results are commonly referred to as binary tree mechanisms. Honaker [20] observed that the full tree used by the binary tree mechanism can be used to produce more efficient estimates since each prefix query can be computed using different combinations of vertices. A subset of Honaker’s techniques can be used to reduce the mean squared error by up to a factor 2, at some cost in efficiency. Rather than leveraging the density of the tree to improve the error, Andersson and Pagh [1] identified that making the tree sparser – only using a subset of the leaves to store inputs – allowed for a factor 2 reduction in the error while keeping the variance of outputs constant, for $(\varepsilon,\delta)$-DP. Although they do not state a result for it, using their technique for $\varepsilon$-DP yields a factor 4 improvement over the binary tree mechanism, which is still more than a factor 2 away from our improvement. We also mention the work of Qardaji, Yang, and Li [25] which, while focusing on range queries, did investigate reducing the mean squared error by means of considering trees of higher arity. They claim an optimal arity of $k=16$, but average over a different set of queries, and do not leverage the subtraction of vertices. Therefore, we emphasize that $k$-ary trees for range queries have already been studied. However, to the best of our knowledge, the connection between $k$-ary numbers and the variance of outputs have not been explicitly made, nor does there seem to be any work on $k$-ary trees for prefix sums in particular.
A different line of work has focused not on trees, but rather attacking the problem from the perspective of factorizing the corresponding query matrix $A\in\{0,1\}^{T\times T}$ — a lower-triangular matrix with all 1s below the diagonal. All known, unbiased, mechanisms for continual counting with good error can be viewed as instantiations in this framework (including our contribution), known as factorization mechanisms or matrix mechanisms [23]. In short, this framework is valid for any set of $q$ linear queries $A\in\mathbb{R}^{q\times n}$ on a dataset $\mathbf{x}\in\mathbb{R}^{n}$ and involves privately releasing $R\mathbf{x}$ by adding noise $\mathbf{z}$ scaled to the sensitivity of $R$, and then outputting $L(R\mathbf{x}+\mathbf{z})$ as post-processing, where the only restriction on $L$ and $R$ is that $LR=A$. Releasing $R\mathbf{x}$ with the Laplacian or Gaussian mechanism yields an $\varepsilon$-DP or $(\varepsilon,\delta)$-DP mechanism respectively.
@@
Multiple recent papers [16, 11, 19] have focused on improving the error for continual counting for $(\varepsilon,\delta)$-DP by directly searching among these factorizations of $A$. Most notably, Henzinger, Upadhyay, and Upadhyay [19] showed that there exists an optimal factorization $L=R=\sqrt{A}$ (also identified by Bennett [3]) whose error is optimal up to the leading constant across all factorizations. But, importantly, this is for $(\varepsilon,\delta)$-DP and assumes using Gaussian noise. Due to the high $\ell_{1}$ sensitivity of $R=\sqrt{A}$, the noise needed to release $R\mathbf{x}$ with the Laplacian mechanism results in a $\varepsilon$-DP matrix mechanism with worse error than that of the binary mechanism.
@@ 2 Preliminaries
Before introducing any mechanism, we first introduce some notation and background. We use $\mathbb{N}$ to denote all positive integers, and for $a,b\in\mathbb{N}$ let $[a,b]\subset\mathbb{N}$ refer to the set of integers $\{a,a+1\dots,b\}$. A vector-valued quantity $\mathbf{y}\in\mathbb{R}^{n}$ is expressed in bold notation and its scalar entries are written as $\mathbf{y}_{i}$ for $i\in[1,n]$. Standard definitions for $\varepsilon$-DP and $(\varepsilon,\delta)$-DP are given in Section A.1.
@@ 3 Mechanisms
In this section, we will gradually build up towards the algorithm outlined in Theorem 1.1. We begin by describing the binary tree mechanism, which will be our starting point. Then we will consider how the utility improves when we extend to $k$-ary trees, and we give an algorithm and results for this case. Finally, our main contributions will be to incorporate subtraction of vertices in $k$-ary trees to further improve the utility. The formal proofs leading to Theorem 1.1 are given in the next section.

### v2

# Count on Your Elders: Laplace vs Gaussian Noise

@@ Abstract
In recent years, Gaussian noise has become a popular tool in differentially private algorithms, often replacing Laplace noise which dominated the early literature on differential privacy. Gaussian noise is the standard approach to approximate differential privacy, often resulting in much higher utility than traditional (pure) differential privacy mechanisms. In this paper we argue that Laplace noise may in fact be preferable to Gaussian noise in many settings, in particular when we seek to achieve $(\varepsilon,\delta)$-differential privacy for small values of $\delta$. We consider two scenarios:
First, we consider the problem of counting under continual observation and present a new generalization of the binary tree mechanism that uses a $k$-ary number system with negative digits to improve the privacy-accuracy trade-off. Our mechanism uses Laplace noise and improves the mean squared error over all “optimal” $(\varepsilon,\delta)$-differentially private factorization mechanisms based on Gaussian noise whenever $\delta$ is sufficiently small. Specifically, using $k=19$ we get an asymptotic improvement over the bound given in the work by Henzinger, Upadhyay and Upadhyay (SODA 2023) when $\delta=O(T^{-0.92})$.
Second, we show that the noise added by the Gaussian mechanism can always be replaced by Laplace noise of comparable variance for the same $(\varepsilon,\delta)$ privacy guarantee, and in fact for sufficiently small $\delta$ the variance of the Laplace noise becomes strictly better. This challenges the conventional wisdom that Gaussian noise should be used for high-dimensional noise.
@@ 1 Introduction
In private data analysis the goal is to release the result of computations on datasets while safeguarding the privacy of the individuals whose data make up the dataset. A popular framework for achieving this is differential privacy [15], which gives probabilistic guarantees on how much can be learned about an individual’s data from the result of a computation. Algorithms providing differential privacy (DP) guarantees conceptually introduce noise, randomly perturbing the result of a corresponding exact, non-private computation. The trade-off between utility and privacy guarantees depends on the nature of the computation.
Let $\mathbf{x},\mathbf{x}^{\prime}\in\{0,1\}^{T}$ be input datasets. We say $\mathbf{x}$ and $\mathbf{x}^{\prime}$ are neighboring if they are equal except for a single bit. We want to compute all prefix sums of $\mathbf{x}$ with differential privacy under this neighboring relation, i.e., we want to privately release $A\mathbf{x}$ where $A$ is the lower-triangular all 1s matrix. If we add the restriction that the bits $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ are received one by one as a stream, and we are to output $(A\mathbf{x})_{i}$ on receiving $\mathbf{x}_{i}$, then this problem is referred to as counting under continual observation or continual counting for short [16, 6].
The standard solution to this problem is the factorization mechanism [30], and it involves (1) factorizing $A$ into matrices $L,R$ such that $A=LR$, (2) privately releasing $R\mathbf{x}$ by adding noise $\mathbf{z}$ and then (3) multiplying the private release by $L$ from the left: $L(R\mathbf{x}+\mathbf{z})=A\mathbf{x}+L\mathbf{z}$. Defining the $\ell_{p}$-sensitivity as $\Delta_{p}=\max_{i\in[T]}\left\lVert R\mathbf{e}_{i}\right\rVert_{p}$, where $\mathbf{e}_{i}$ is the $i$th unit basis vector, classical results on DP say that adding Laplace noise scaled to $\Delta_{1}$ yields $\varepsilon$-DP (also called pure DP), and adding Gaussian noise scaled to $\Delta_{2}$ yields $(\varepsilon,\delta)$-DP (also called approximate DP). The scheme rests on the observation that while the sensitivity of $A$ may be large, the sensitivity of $R$ can be much smaller, and so adding noise $L\mathbf{z}$ to the final result may result in better accuracy than adding noise directly to $A\mathbf{x}$.
While continual counting is a deceptively simple problem, we note that it still has signifcant gaps between upper and lower bounds [16, 24, 19, 17, 9]. The problem is used as a subroutine in various applications ranging from private learning [32, 28, 8, 12, 7] to histogram estimation [4, 5, 27, 36], so any improvement in upper bounds affect these applications.
@@ 1.1 Our contributions
Our first contribution is an easy-to-implement tree-aggregation-based mechanism for continual counting under $\varepsilon$-DP, and based on Laplace noise, with properties stated in Theorem 1.1.
@@
Here and elsewhere $\log_{k}$ is the base $k$ logarithm, and $\log(\cdot)=\log_{2}(\cdot)$. To the best of our knowledge, this is the best bound on the mean squared error known for any $\varepsilon$-DP mechanism, up to lower-order asymptotically vanishing terms. It improves the error achieved by applying techniques in [26] to the binary tree mechanism (as implemented in [28] and referred to elsewhere as ?Honaker Online?) by a factor 4 when $k=19$. While this is interesting in its own right, we note that this improvement in the leading constant has implications for how the error compares to $(\varepsilon,\delta)$-DP mechanisms. Henzinger, Upadhyay and Upadhyay [24] produce a factorization with a bound on the mean squared error of $C_{\varepsilon,\delta}^{2}\big{(}1+\frac{\ln(4T/5)}{\pi}\big{)}^{2}$, where $C_{\varepsilon,\delta}=\frac{2}{\varepsilon}\sqrt{\frac{4}{9}+\ln\big{(}\frac{1}{\delta}\sqrt{\frac{2}{\pi}}\big{)}}$, and show that no factorization mechanism based on Gaussian noise can asymptotically improve on this error. [^1] As the constant $C_{\varepsilon,\delta}$ is valid for all $\varepsilon,\delta\in(0,1)$, it seems plausible that it can be reduced via a tighter analysis of the Gaussian mechanism, as done in for example [2]. We improve on their error using Laplacian noise for small enough $\delta$.
Theorem 1.2.
Given an integer $T\geq 2$, there exists an $\varepsilon$-DP algorithm for continual counting, based on Laplacian noise, that improves on the mean squared error bound given in [24], up to lower-order asymptotically vanishing terms, whenever $\delta=O(T^{-0.92})$.
Given that $\delta$ is commonly set to be $o(1/T)$, this result is surprising. The phenomenon is discussed in Section 5, but in a nutshell a sufficiently small $\delta$ must cause the error to greatly increase if the mechanism is based on Gaussian noise. While improving constants for $\varepsilon$-DP is of independent interest, Theorem 1.2 implies that further improvements also have implications for $(\varepsilon,\delta)$-DP.
Motivated by the dominance of Gaussian noise in applications, and the moral in Theorem 1.2 for Laplace noise, our second contribution is $(\varepsilon,\delta)$-DP guarantees for the Laplace mechanism for given $\ell_{1}$ and $\ell_{2}$ sensitivities.
Theorem 1.3 (Approximate DP for the Laplace Mechanism).
Let $f:\mathcal{X}^{n}\to\mathbb{R}^{d}$ be a function with $\ell_{p}$-sensitivity $\Delta_{p}\coloneqq\max_{\mathcal{D}\sim\mathcal{D}^{\prime}}\left\lVert f(\mathcal{D})-f(\mathcal{D}^{\prime})\right\rVert_{p}$. For a given dataset $\mathcal{D}\in\mathcal{X}^{n}$, $\delta\in(0,1)$ and $\lambda>\Delta_{1}$, the output $f(\mathcal{D})+\mathsf{Lap}(\lambda)^{d}$, satisfies $(\varepsilon,\delta)$-differential privacy where
$\varepsilon=\min\bigg{\{}\frac{\Delta_{1}}{\lambda},\frac{\Delta_{2}}{\lambda}\bigg{(}\frac{\Delta_{2}}{2\lambda}+\sqrt{2\ln(1/\delta)}\bigg{)}\bigg{\}}\,.$
Essentially we can implement $(\varepsilon,\delta)$-DP for the Laplace mechanism where the resulting $\varepsilon$ is never worse than what simple composition gives, while also being competitive with the Gaussian mechanism in the regime where it is advantageous.
@@ 1.2 Overview of technical ideas
We start out by building a complete $k$-ary tree of height $h=\lceil\log_{k}(T+1)\rceil$ on top of the input $\mathbf{x}\in\{0,1\}^{T}$, see Figure 1. To analyze this, it turns out that expressing a time step $t$ as a $k$-ary number, i.e., as $h$ digits in $\{0,\dots,k-1\}$, i.e. $\mathbf{t}\in\{0,\dots,k-1\}^{h}$, gives a direct means of computing which vertices contribute to a given prefix sum. Increasing the arity of the tree reduces the height, which reduces the $\ell_{1}$-sensitivity and levels of vertices that might get added, but increases the number of vertices to add per level. To mitigate the increase in number of vertices to add, we leverage subtraction, see Figure 2
To analyze the structure of the resulting prefix sums when subtraction is introduced, it turns out that rather than expressing $t$ in digits from 00 to $k-1$, we should instead allow for negative digits. Focusing on the case of odd $k$, we thus represent $t$ as $\mathbf{t}\in\{-\frac{k-1}{2},\dots,\frac{k-1}{2}\}^{h}$. The interpretation of $\mathbf{t}$ becomes: if $\mathbf{t}_{\ell}\geq 0$, add $\mathbf{t}_{\ell}$ leftmost children on level $\ell$, otherwise, subtract $|\mathbf{t}_{\ell}|$ rightmost children on level $\ell$. Identifying that $\left\lVert\mathbf{t}\right\rVert_{1}$ is equal to the number of vertices used to produce the estimate at time $t$, we can analyze the average number of vertices added for an estimate from this representation by studying $\mathbf{t}$ for $t\in[1,T]$. This gives, roughly, a factor-2 reduction in the mean squared error versus not using subtraction.
@@
To derive the $(\varepsilon,\delta)$-DP guarantee of the Laplace mechanism, we make use of a heterogeneous composition theorem [29]. Given two arbitrary multidimensional neighboring inputs, we argue for the privacy loss in each coordinate when one of them is released with respect to its neighbor. To get the aggregate privacy guarantee for the entire vector, we apply the composition theorem over the individual coordinates being released, and show that the guarantee must hold for any two neighboring outputs. It turns out that, as in the case of the Gaussian mechanism, the $(\varepsilon,\delta)$-DP guarantee for Laplacian noise naturally depends on the $\ell_{2}$-sensitivity of the function being released.
@@ 1.3 Related work
Historically, the first mechanisms with good error are typicallywcredited to Chan, Shi, and Song [6] and Dwork, Naor, Pitassi, and Rothblum [16], though similar constructions were proposed independently by others [23, 20]. These results are commonly referred to as binary tree mechanisms. Honaker [26] observed that the full tree used by the binary tree mechanism can be used to produce more efficient estimates since each prefix query can be computed using different combinations of vertices. A subset of Honaker’s techniques can be used to reduce the mean squared error by up to a factor 2, at some cost in efficiency. Rather than leveraging the density of the tree to improve the error, Andersson and Pagh [1] identified that making the tree sparser – only using a subset of the leaves to store inputs – allowed for a factor 2 reduction in the error while keeping the variance of outputs constant, for $(\varepsilon,\delta)$-DP. Although they do not state a result for it, using their technique for $\varepsilon$-DP yields a factor 4 improvement over the binary tree mechanism, which is still more than a factor 2 away from our improvement. We also mention the work of Qardaji, Yang, and Li [33] which, while focusing on range queries, did investigate reducing the mean squared error by means of considering trees of higher arity. They claim an optimal arity of $k=16$, but average over a different set of queries, and do not leverage the subtraction of vertices. Therefore, we emphasize that $k$-ary trees for range queries have already been studied. However, to the best of our knowledge, the connection between $k$-ary numbers and the variance of outputs have not been explicitly made, nor does there seem to be any work on $k$-ary trees for prefix sums in particular.
@@
Many recent papers [14, 12, 24, 19] have focused on improving the error for continual counting for $(\varepsilon,\delta)$-DP by directly searching among factorizations of $A$, the lower-triangular all-1s matrix. Most notably, Henzinger, Upadhyay, and Upadhyay [24] showed that there exists an optimal factorization $L=R=\sqrt{A}$ (also identified by Bennett [3]) whose error is optimal up to the leading constant across all factorizations. But, importantly, this is for $(\varepsilon,\delta)$-DP and assumes using Gaussian noise. Due to the high $\ell_{1}$-sensitivity of $R=\sqrt{A}$, the noise needed to release $R\mathbf{x}$ with the (conventional) Laplacian mechanism results in an $\varepsilon$-DP matrix mechanism with worse error than that of the binary mechanism.
The second part of our paper explores the use of Laplace noise for $(\varepsilon,\delta)$-DP, especially in the small $\delta$ regime – a regime we are not the first to investigate [35]. We claim no novelty in pointing out that the Gaussian mechanism scales poorly for small $\delta$ (see e.g. [13]). We do note that there is a rich literature investigating alternative (non-Gaussian) noise distribution for approximate DP [22, 25, 31, 11, 21, 37]. In [34] they compare how Laplace and Gaussian noise behaves over many compositions, and especially observe that eventually they will both converge to Gaussian privacy loss distributions.
@@ 2 Preliminaries
Before introducing any mechanism, we first introduce some notation and background. We use $\mathbb{N}$ to denote all positive integers, and for $a,b\in\mathbb{N}$ let $[a,b]\subset\mathbb{N}$ refer to the set of integers $\{a,a+1\dots,b\}$. A vector-valued quantity $\mathbf{y}\in\mathbb{R}^{n}$ is expressed in bold notation and its scalar entries are written as $\mathbf{y}_{i}$ for $i\in[1,n]$. Standard definitions for $\varepsilon$-DP and $(\varepsilon,\delta)$-DP are given in Appendix A.
@@ 3 Mechanisms
In this section, we will gradually build up towards the algorithm outlined in Theorem 1.1. We begin by describing the binary tree mechanism, which will be our starting point. Then we will consider how the utility improves when we extend to $k$-ary trees, and we give an algorithm and results for this case. Finally, our main contribution will be to incorporate subtraction of vertices in $k$-ary trees to further improve the utility. The formal proofs leading to Theorem 1.1 are given in the next section.